OpenStudy (anonymous):

A cell and its nucleus both have a spherical shape; the nucleus is 1 µm diameter and the cell is 15 um in diameter. ( Note: the formula for the surface area is 4πr2, and the formula for the volume of a sphere is : 4/3πr3, where r is the radius of the sphere and the π is 3.1416.) After calculating the cellular/ nuclear volume, how many times greater would the cell/nuclear ratio be if the cell were to triple its diameter (assuming the nucleus would remain the same size)?

OpenStudy (anonymous):

just put the values in the formula and calculate it u will find the ratio

OpenStudy (aaronq):

V= (4/3)*π*r^3=(4/3)*π*(d/2)^3 so first calculate the volume of both as they are: d=1 micrometer and 10 micrometers. (use: V=(4/3)*π*(d/2)^3) -> find the ratio, V(nucleus)/V(cell) then, if the cell triples diameter, so 10*3 =30 micrometers. find it's volume, again find the ratio, V(nucleus)/V(cell), with this new value. now divide the second ratio by the first ratio. second ratio/first ratio= how many time greater the ratio would be if d of cell was x3.